# Analysis of Fourier Neural Operators via Effective Field Theory

### v2

@@ Abstract
Fourier Neural Operators (FNOs) have emerged as leading surrogates for solver operators for various functional problems, yet their stability, generalization and frequency behavior lack a principled explanation. We present a systematic effective field theory analysis of FNOs in an infinite dimensional function space, deriving closed recursion relations for the layer kernel and four point vertex and then examining three practically important settings—analytic activations, scale invariant cases and architectures with residual connections. The theory shows that nonlinear activations inevitably couple frequency inputs to high frequency modes that are otherwise discarded by spectral truncation, and experiments confirm this frequency transfer. For wide networks, we derive explicit criticality conditions on the weight initialization ensemble that ensure small input perturbations maintain a uniform scale across depth, and we confirm experimentally that the theoretically predicted ratio of kernel perturbations matches the measurements. Taken together, our results quantify how nonlinearity enables neural operators to capture non-trivial features, supply criteria for hyperparameter selection via criticality analysis, and explain why scale invariant activations and residual connections enhance feature learning in FNOs.
@@ 1.1 Fourier Neural Operator
In scientific machine learning, neural operators—networks that approximate solution operators so they can act directly on functional data—are attracting growing attention (Kovachki et al. (2021b), Lu et al. (2021)). In addition to the linear layers of standard fully connected networks (FCNs), neural operator architectures include layers that approximate a kernel and apply an integral transform. How the kernel is represented gives rise to many variants: graph based constructions (Li et al. (2020)), tensor product decompositions (Kovachki et al. (2021b)), hierarchical (multi-resolution) kernels (Gupta et al. (2021)), and, most notably, Fourier Neural Operators (FNOs) (Li et al. (2021)). Thanks to their computational efficiency, high accuracy, and ability to capture long range interactions that elude CNNs and GNNs, FNOs have become the workhorse choice for surrogate modeling (Pathak et al. (2022), Sun et al. (2023)). Prior theory has addressed their generalization error (Kim & Kang (2024), Benitez et al. (2024)), the universal approximation property (Kovachki et al. (2021a), Lee et al. (2025)), and expressivity/trainability from a mean-field perspective (Koshizuka et al. (2024)), yet a comprehensive statistical description is still lacking. Here we fill that gap by analyzing FNOs through the lens of effective field theory (EFT).
@@ 1.3 Contributions
Building on the EFT formulation for FCNs, we extend the approach to Fourier Neural Operators—an essential step, because FNOs act on infinite dimensional function spaces rather than finite-dimensional vectors. While the basic relations among connected correlators mirror those in FCNs, the architectural differences introduce qualitatively new phenomena: the observables become functions rather than scalars, and mechanisms such as frequency coupling appear. We show that the kernel function—the infinite width approximation of the two-point correlator—admits a closed layerwise recursion, and we derive analytic formulas for kernel perturbations; we then validate these predictions experimentally. Our analysis quantifies how non-trivial activations and residual connections influence information flow in FNOs and yields explicit design criteria for stable models.
2 Preliminary
@@ 2.2 Neural Operator
The architecture of the neural operator we consider consists of the following two types of layers: Eq. 1 describes a linear layer that linearly transforms the channel dimension of functional data, and Eq. 2 represents a kernel integration layer performing convolution operations with a kernel function. Eq. 3 expresses the kernel integration layer combined with a non-linear activation function. In general, the kernel function does not need to be translationally invariant; however, in our setting, we focus on the translationally invariant case as expressed in Eq. 2.
@@
$\Big{(}\mathcal{R}(u)(x)\Big{)}_{i}:=\int_{\mathbb{R}^{d}}k_{ij}(x-x^{\prime})u_{j}(x^{\prime})dx^{\prime},\quad i=1,\dots,n.$ (2)
@@
$\begin{split}\mathcal{Z}^{(1)}&:=\mathcal{L}_{\text{lift}}(u)\\ \mathcal{Z}^{(l+1)}(u)&:=\mathcal{R}^{(l+1)}\Big{(}\mathfrak{S}^{(l)}(u)\Big{)},\quad l=1,\dots,L-1,\\ \mathcal{Y}(u)&:=\mathcal{L}_{\text{proj}}\Big{(}\mathcal{Z}^{(L)}(u)\Big{)}.\end{split}$ (4)
@@ 3 Effective Field Theory for Neural Operators
Because both differentiation and ordinary integration act linearly on the functional measure, connected correlators furnish a systematic means of computing expectation values that involve products, derivatives, or integrals of multiple fields. In what follows, closed form expressions for these correlators are derived, and—paralleling the analysis in Appendix A—recursive relations are established for the two-point kernel and the four-point vertex. Finally, explicit formulas are provided for the susceptibility of the kernel to perturbations parallel and perpendicular to a chosen reference trajectory in function space, thereby quantifying anisotropic responses to input variations.
@@ 3.1 Correlators for Neural Operators
Henceforth, boldface letters indicate vector-valued functions. And notation $\hat{u}$ also means $\mathcal{F}(u)$. Because the explicit evaluation of correlators in layers that include convolution is prohibitively cumbersome, we instead work in Fourier space and compute the correlators of the Fourier transformed pre-activations. In particular, Eq. 6 gives the two-point correlator of the $l$-th layer conditioned on the (l1)(l-1)-th layer’s outputs $\mathbf{u}$ and $\mathbf{v}$. We write $\Big{(}\mathcal{Z}^{(l)}|_{\mathbf{u}}\Big{)}_{i}$ for the $i$-th component of $\mathcal{Z}^{(l)}|_{\mathbf{u}}$, that is, the $i$-th pre-activation in layer $l$ when the preceding layer is fixed to the function $\mathbf{u}$. Likewise, we write $\Big{(}\mathcal{Z}^{(l)}\{\mathbf{u}\}\Big{)}_{i}$ for the corresponding output component of layer $l$ when the input function is $\mathbf{u}$.
@@
$\begin{split}&\mathbb{E}\bigg{[}\Big{(}\mathcal{Z}^{(l)}|_{\mathbf{u}}(x)\Big{)}_{i}\overline{\Big{(}\mathcal{Z}^{(l)}|_{\mathbf{v}}(y)\Big{)}}_{i^{\prime}}\bigg{|}\mathbf{u},\mathbf{v}\bigg{]}\\ &=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}\mathcal{F}\Big{(}\mathcal{Z}^{(l)}|_{\mathbf{u}}\Big{)}_{i}(f)\overline{\mathcal{F}\Big{(}\mathcal{Z}^{(l)}|_{\mathbf{v}}\Big{)}}_{i^{\prime}}(f^{\prime})\bigg{|}\mathbf{u},\mathbf{v}\bigg{]}e^{ifx}e^{-if^{\prime}y}dfdf^{\prime}\\ &=\frac{C_{R^{(l)}}}{n^{(l-1)}}\delta_{ii^{\prime}}\mathcal{F}^{-1}\Big{(}\langle\mathbf{\hat{u}},\mathbf{\hat{v}}\rangle\Big{)}\end{split}$ (7)
Now, we consider two-point correlator for neural operator, since $l$-th layer parameters and variables up to (l1)(l-1)-th layers are independent, the two-point correlator of a neural operator factorizes as Eq. 8.
$\begin{split}&\mathbb{E}\bigg{[}\mathcal{F}\Big{(}\mathcal{Z}^{(l)}\{\mathbf{u}\}\Big{)}_{i}(f)\overline{\mathcal{F}\Big{(}\mathcal{Z}^{(l)}\{\mathbf{v}\}\Big{)}_{i^{\prime}}}(f^{\prime})\bigg{]}\\ &=\frac{C_{R^{(l)}}}{n^{(l-1)}}\delta_{ii^{\prime}}\delta(f-f^{\prime})\mathbb{E}\bigg{[}\Big{(}\mathcal{F}\Big{(}\mathfrak{S}^{(l-1)}\{\mathbf{u}\}\Big{)}(f)\cdot\overline{\mathcal{F}\Big{(}\mathfrak{S}^{(l-1)}\{\mathbf{v}\}\Big{)}}(f^{\prime})\bigg{]}\end{split}$ (8)
and we define the $l$-th layer mean metric $\mathcal{G}^{(l)}$ as non-trivial part of Eq. 8 as Eq. 9 and stochastic metric $\widetilde{\mathcal{G}^{(l)}}$ as Eq. 10 in which expectation is not taken except on last $l$-th layer:
$\mathcal{G}^{(l)}\{\mathbf{u},\mathbf{v}\}(f,f^{\prime}):=\frac{C_{R^{(l)}}}{n^{(l-1)}}\mathbb{E}\bigg{[}\mathcal{F}\Big{(}\mathfrak{S}^{(l-1)}\{\mathbf{u}\}\Big{)}(f)\cdot\overline{\mathcal{F}\Big{(}\mathfrak{S}^{(l-1)}\{\mathbf{v}\}\Big{)}}(f^{\prime})\bigg{]}$ (9)
$\widetilde{\mathcal{G}^{(l)}}\{\mathbf{u},\mathbf{v}\}(f,f^{\prime}):=\frac{C_{R^{(l)}}}{n^{(l-1)}}\sum_{j}\mathcal{F}\Big{(}\mathfrak{S}^{(l-1)}\{\mathbf{u}\}\Big{)}_{j}(f)\overline{\mathcal{F}\Big{(}\mathfrak{S}^{(l-1)}\{\mathbf{v}\}\Big{)}_{j}}(f^{\prime})$ (10)
@@ 3.2 Running of Couplings
In this subsection we derive recursion relations for the correlation functions—what, in field theory language, is the running of couplings. As the layer width tends to infinity, the pre-activation ensemble becomes Gaussian and the layer metric converges to a deterministic kernel $\mathcal{K}^{(l)}$ (see Appendix B). In this vanishing fluctuation regime, the next layer kernel can be expressed as a Gaussian expectation with covariance $\mathcal{K}^{(l)}$ inherited from the previous layer; equivalently, $\mathcal{K}^{(l+1)}$ is obtained by integrating over the Gaussian process defined by $\mathcal{K}^{(l)}$.
$\begin{split}\mathcal{K}^{(l+1)}\{\mathbf{u},\mathbf{v}\}(f,f^{\prime}):=C_{R^{(l+1)}}(f)\langle\mathcal{F}(\mathfrak{S}^{(l)}\{\mathbf{u}\}),\mathcal{F}(\mathfrak{S}^{(l)}\{\mathbf{v}\})\rangle_{\mathcal{K}^{(l)}}(f,f^{\prime}).\end{split}$
@@
$\begin{split}\mathcal{K}^{(1)}\{\mathbf{u},\mathbf{v}\}(f,f^{\prime}):=\frac{C_{W}}{n^{(1)}}\sum_{j}\hat{u_{j}}(f)\hat{v_{j}}(f^{\prime})\end{split}$ (11)
$\begin{split}\mathcal{K}^{(l+1)}\{\mathbf{u}\}(f,f^{\prime})=\frac{C_{R^{(l+1)}}}{n^{(l)}}g(\mathcal{K}^{(l)}\{\mathbf{u}\}(f,f^{\prime}))\\ g(\mathcal{K}):=\langle\mathcal{F}(\mathfrak{S}^{(l)}\{\mathbf{u}\}),\mathcal{F}(\mathfrak{S}^{(l)}\{\mathbf{u}\})\rangle_{\mathcal{K}^{(l)}}(f,f^{\prime}).\end{split}$ (12)
@@
$\tilde{\chi}_{\perp}(f,f^{\prime}):=\|\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\})\|_{\mathcal{K}^{(l)}\{u_{0}\}}$
@@
Here, $\tilde{\chi}_{\perp}$ is calculated at layer $l$ and $C(f)$ is $C_{R^{(l+1)}}$.
4 Special Cases
In this section we put the general framework of Section 3 to work by carrying out the calculation in three representative settings—(i) analytic activations, (ii) scale invariant architectures, and (iii) networks equipped with residual connections.

### v3

@@ Abstract
Fourier Neural Operators (FNOs) have emerged as leading surrogates for solver operators for various functional problems, yet their stability, generalization and frequency behavior lack a principled explanation. We present a systematic effective field theory analysis of FNOs in an infinite-dimensional function space, deriving closed recursion relations for the layer kernel and four-point vertex and then examining three practically important settings—analytic activations, scale-invariant cases and architectures with residual connections. The theory shows that nonlinear activations inevitably couple frequency inputs to high frequency modes that are otherwise discarded by spectral truncation, and experiments confirm this frequency transfer. For wide networks, we derive explicit criticality conditions on the weight initialization ensemble that ensure small input perturbations maintain a uniform scale across depth, and we confirm experimentally that the theoretically predicted ratio of kernel perturbations matches the measurements. Taken together, our results quantify how nonlinearity enables neural operators to capture non-trivial features, supply criteria for hyperparameter selection via criticality analysis, and explain why scale-invariant activations and residual connections enhance feature learning in FNOs. Finally, we translate the criticality theory into a practical criterion-matched initialization (calibration) procedure; on a standard PDEBench Burgers benchmark, the calibrated FNO exhibits markedly more stable optimization, faster convergence, and improved test error relative to a vanilla FNO.
Keywords: Fourier neural operator; effective field theory; criticality condition; kernel recursion
@@ 1.1 Fourier Neural Operator
In scientific machine learning, neural operators—networks that approximate solution operators so they can act directly on functional data—are attracting growing attention (Kovachki et al. (2021b), Lu et al. (2021)). In addition to the linear layers of standard fully connected networks (FCNs), neural operator architectures include layers that approximate a kernel and apply an integral transform. How the kernel is represented gives rise to many variants: graph-based constructions (Li et al. (2020)), tensor product decompositions (Kovachki et al. (2021b)), hierarchical (multi-resolution) kernels (Gupta et al. (2021)), and, most notably, Fourier Neural Operators (FNOs) (Li et al. (2021)). Thanks to their computational efficiency, high accuracy, and ability to capture long-range interactions that elude CNNs and GNNs, FNOs have become the workhorse choice for surrogate modeling (Pathak et al. (2022), Sun et al. (2023)). Prior theory has addressed their generalization error (Kim and Kang (2024a), Benitez et al. (2024)), the universal approximation property (Kovachki et al. (2021a), Lee et al. (2025)), and expressivity/trainability from a mean-field perspective (Koshizuka et al. (2024)), yet a comprehensive statistical description is still lacking. Here we fill that gap by analyzing FNOs through the lens of effective field theory (EFT).
@@ 1.3 Contributions
Building on the EFT formulation for FCNs, we extend the approach to Fourier Neural Operators—an essential step, because FNOs act on infinite dimensional function spaces rather than finite-dimensional vectors. While the basic relations among connected correlators mirror those in FCNs, the architectural differences introduce qualitatively new phenomena: the observables become functions rather than scalars, and mechanisms such as frequency coupling appear. While frequency-coupling phenomena can be inferred qualitatively, our EFT framework explains the quantitative decay of coupling strength and its dependence on network width. We show that the kernel function—the infinite width approximation of the two-point correlator—admits a closed layerwise recursion, and we derive analytic formulas for kernel perturbations; we then validate these predictions experimentally. Our analysis quantifies how non-trivial activations and residual connections influence information flow in FNOs and yields explicit design criteria for stable models. Finally, leveraging these theoretical insights, we propose a practical criterion-matched initialization (calibration) algorithm; on a real benchmark dataset, it leads to more stable training dynamics, faster convergence, lower achieved loss, and improved generalization compared to the vanilla FNO baseline.
Fourier Neural Operator (FNO) with spectral truncation.  Given an input function $u(x)$, each Fourier layer applies $\mathrm{FFT}\rightarrow$ truncation ($|f|\leq k_{\max}$) $\rightarrow$ multiplication by $R^{(l)}(f)$ $\rightarrow\mathrm{iFFT}$, followed by a pointwise nonlinearity $\sigma$; optionally, a residual update $Z^{(l+1)}=\mathcal{R}^{(l+1)}(\sigma(\cdot))+\tilde{\gamma}Z^{(l)}$ is used to improve depth-wise information retention. Mechanism of frequency transfer.  A pointwise nonlinearity becomes a convolutional expansion in Fourier space, $\mathcal{F}(\sigma(g))(f)=\sum_{n=1}^{\infty}\frac{\sigma_{n}}{n!}\hat{g}^{*n}(f).$ so nonlinear layers inevitably generate modes beyond the truncation band. For analytic activations, the contribution of the $k$-th term is suppressed by the width as $\mathcal{O}(n^{-(k-1)})$. EFT formulation in function space.  We treat random initialization as an ensemble over parameters (with fixed input functions $u,v$) and track connected correlators of the Fourier-domain pre-activations. The principal observables are the two-point kernel $\mathcal{K}^{(l)}$ and the four-point vertex $\mathcal{V}^{(l)}$. In the infinite-width limit, pre-activations converge to a Gaussian process and $\mathcal{V}^{(l)}=\mathcal{O}(1/n)$, yielding closed layerwise recursions for $\mathcal{K}^{(l+1)}$ and for the parallel/perpendicular susceptibilities $\chi_{\parallel}$ and $\chi_{\perp}$, which determine criticality conditions for stable signal propagation. Three representative regimes.  (i) Analytic activations: explicit convolution-series recursions for $\mathcal{K}^{(l)}$ and susceptibilities. (ii) Scale-invariant (ReLU-type) activations: kernels are governed by the position-space correlation $\rho^{(l)}$, leading to broad spectral support and slow decay. (iii) Residual FNOs: residual gain $\gamma$ controls the retention of post-truncation energy and enables deeper compositions while preserving truncation efficiency. Empirical validation.  Across activation families (tanh/polynomial/ReLU) and with/without residual connections, the kernel recursion accurately predicts the measured kernel evolution, including energy in post-truncation frequencies. Measured susceptibilities $\chi_{\parallel}$ and $\chi_{\perp}$ agree with theory within one-standard-deviation bands over repeated random initializations. Importantly, applying the proposed calibration at initialization translates to improved downstream training performance, yielding more stable optimization and consistently better final metrics (e.g., faster convergence and/or lower test error) than uncalibrated baselines.
Figure 1: Schematic overview of the paper: (top) the truncated FNO architecture; (middle) nonlinearity-induced frequency transfer and the EFT description via the kernel/vertex hierarchy; (bottom) closed-form results in three regimes and their experimental validation and the effectiveness of calibration based on our theory.
2 Preliminaries
@@ 2.2 Neural Operator
The architecture of the neural operator we consider consists of the following two types of layers: Eq. 1 describes a linear layer that linearly transforms the channel dimension of functional data, and Eq. 2 represents a kernel integration layer performing convolution operations with a kernel function. Eq. 3 expresses the kernel integration layer combined with a non-linear activation function. In general, the kernel function does not need to be translationally invariant; however, in our setting, we focus on the translationally invariant case as explained in assumption (A1).
@@
$\Big(\mathcal{R}(u)(x)\Big)_{i}:=\int_{\mathbb{R}^{d}}k_{ij}(x,x^{\prime})u_{j}(x^{\prime})dx^{\prime},\quad i=1,\dots,n.$ (2)
@@
Modeling assumptions and scope.
For clarity, we list the assumptions under which the main EFT recursions and closed-form results in Sections 3–4 are derived.
(A1) Translation-invariant kernel layers. Each kernel integration layer is a translation-invariant convolution operator $k(x-x^{\prime})$, so that in Fourier space the layer acts by mode-wise multiplication.
(A2) Channel-i.i.d. Gaussian initialization. For each Fourier mode, the complex-valued parameters $\{R^{(l)}_{ij}(f)\}_{i,j}$ are initialized as mean-zero Gaussian variables, independent across channel indices $(i,j)$ and across layers, with variance scaling $1/n^{(l-1)}$.
(A3) Frequency-diagonal spectral covariance (baseline setting). In the main development we assume frequency-wise independence, $\mathbb{E}[R^{(l)}_{ij}(f)\overline{R^{(l)}_{i^{\prime}j^{\prime}}(f^{\prime})}]\propto\delta(f-f^{\prime})$, which yields the diagonal spectral profile $C_{R^{(l)}}(f)\delta(f-f^{\prime})$ in Eq. 5. Truncation is encoded by choosing $C_{R^{(l)}}(f)=0$ for $|f|>k_{\max}$.
Covered architectures. Assumptions (A1)–(A3) include the standard FNO layer (FFT $\rightarrow$ truncation $\rightarrow$ multiplication by a learnable tensor $\rightarrow$ iFFT $\rightarrow\sigma$), arbitrary fixed spectral profiles $C_{R}(f)$ (step, decaying, etc.), and ResNet-style variants as in Eq. 16. Linear lift/projection layers initialized as in the FCN baseline are also covered. Moreover, the CNN subnetwork commonly used in standard FNO implementations can be interpreted in our framework as a Fourier layer whose effective kernel is spatially localized (equivalently, whose spectral response is broadly supported rather than concentrated on low modes); hence, its contribution can be analyzed within the same operator-kernel formalism.
Outside the present scope. Architectures that break translation invariance at the kernel level (e.g., explicitly $x$-dependent kernels, windowed/local operators that are not diagonalizable by a global Fourier basis), structured parameterizations that couple channels and frequencies beyond the Gaussian ensemble above, or non-Gaussian initializations may require modified correlator calculations. We discuss a direct generalization to correlated Gaussian spectral weights below.
@@
$\begin{split}\mathcal{Z}^{(1)}&:=\mathcal{L}_{\text{lift}}(u),\\ \mathcal{Z}^{(l+1)}(u)&:=\mathcal{R}^{(l+1)}\Big(\mathfrak{S}^{(l)}(u)\Big),\quad l=1,\dots,L-1,\\ \mathcal{Y}(u)&:=\mathcal{L}_{\text{proj}}\Big(\mathcal{Z}^{(L)}(u)\Big).\end{split}$ (4)
@@
Remark (Lemma 2 and the frequency-diagonal assumption).
Lemma 2 relies critically on the frequency-diagonal covariance in Assumption (A3), $\mathbb{E}[R(f)\overline{R(f^{\prime})}]=C(f)\delta(f-f^{\prime}).$ This $\delta$-structure forces the Wick contractions to collapse onto a single frequency variable, yielding the closed one-dimensional convolution form $\delta(f-f^{\prime})\,H^{*n}(f)$. If we instead allow correlated Gaussian spectral weights with a non-diagonal covariance $\mathbb{E}[R(f)\overline{R(f^{\prime})}]=\Sigma_{R}(f,f^{\prime}),$ then the contractions no longer produce $\delta(f-f^{\prime})$, and the conclusion of Lemma 2 does not hold in this simple form: the resulting moments remain genuinely two-frequency objects (with sums over pairings/permutations), and the recursion formulas in Theorem 1 must be written in terms of the full kernel $\mathcal{K}^{(l)}(f,f^{\prime})$ rather than its diagonal.
Most practical FNO implementations parameterize Fourier modes independently (up to the conjugate-symmetry constraint for real outputs) and apply spectral truncation mode-wise, which corresponds precisely to Assumption (A3) with a diagonal profile $C_{R}(f)\delta(f-f^{\prime})$. Our experiments follow this standard implementation choice; it preserves the FFT-based efficiency and yields a frequency-diagonal kernel that admits the explicit convolution recursions used throughout the paper. Allowing general $\Sigma_{R}(f,f^{\prime})$ is a meaningful extension, but it typically requires tracking dense frequency–frequency correlations and leads to substantially more involved analytic expressions.
@@ 3 Effective Field Theory for Neural Operators
Because both differentiation and ordinary integration act linearly on the functional measure, connected correlators furnish a systematic means of computing expectation values that involve products, derivatives, or integrals of multiple fields. In what follows, closed form expressions for these correlators are derived, and—paralleling the analysis in Appendix A—recursive relations are established for the two-point kernel and the four-point vertex. Finally, explicit formulas are provided for the susceptibility of the kernel to perturbations parallel and perpendicular to a chosen reference trajectory in function space, thereby quantifying anisotropic responses to input variations. In the analysis, $\mathbf{u}$ and $\mathbf{v}$ are fixed input functions; no assumption on an input distribution is needed. Expectations are taken over the random initialization of the network parameters, while in experiments we use a Gaussian random field for sampling input functions only as a convenient testbed.
@@ 3.1 Correlators for Neural Operators
Henceforth, boldface letters indicate vector-valued functions. The notation $\hat{u}$ also means $\mathcal{F}(u)$. Because the explicit evaluation of correlators in layers that include convolution is prohibitively cumbersome, we instead work in Fourier space and compute the correlators of the Fourier transformed pre-activations. In particular, Eq. 6 gives the two-point correlator of the $l$-th layer conditioned on the (l1)(l-1)-th layer’s outputs $\mathbf{u}$ and $\mathbf{v}$. We write $\Big(\mathcal{Z}^{(l)}|_{\mathbf{u}}\Big)_{i}$ for the $i$-th component of $\mathcal{Z}^{(l)}|_{\mathbf{u}}$, that is, the $i$-th pre-activation in layer $l$ when the preceding layer is fixed to the function $\mathbf{u}$. Likewise, we write $\Big(\mathcal{Z}^{(l)}\{\mathbf{u}\}\Big)_{i}$ for the corresponding output component of layer $l$ when the input function is $\mathbf{u}$.
@@
[…]
Now we consider the two-point correlator for the neural operator. Since the parameters in $l$-th layer are independent of the variables up to (l1)(l-1)-th layers, the two-point correlator factorizes as in Eq. 8.
$\begin{split}&\mathbb{E}\bigg[\mathcal{F}\Big(\mathcal{Z}^{(l)}\{\mathbf{u}\}\Big)_{i}(f)\overline{\mathcal{F}\Big(\mathcal{Z}^{(l)}\{\mathbf{v}\}\Big)_{i^{\prime}}}(f^{\prime})\bigg]\\ &=\frac{C_{R^{(l)}}}{n^{(l-1)}}\delta_{ii^{\prime}}\delta(f-f^{\prime})\mathbb{E}\bigg[\Big(\mathcal{F}\Big(\mathfrak{S}^{(l-1)}\{\mathbf{u}\}\Big)(f)\cdot\overline{\mathcal{F}\Big(\mathfrak{S}^{(l-1)}\{\mathbf{v}\}\Big)}(f^{\prime})\bigg].\end{split}$ (8)
We define the $l$-th layer mean metric $\mathcal{G}^{(l)}$ as the non-trivial part of Eq. 8, i.e., Eq. 9, and define stochastic metric $\widetilde{\mathcal{G}^{(l)}}$ as in Eq. 10, where the expectation is taken only over the parameters of $l$-th layer:
$\mathcal{G}^{(l)}\{\mathbf{u},\mathbf{v}\}(f,f^{\prime}):=\delta(f-f^{\prime})\frac{C_{R^{(l)}}}{n^{(l-1)}}\mathbb{E}\bigg[\mathcal{F}\Big(\mathfrak{S}^{(l-1)}\{\mathbf{u}\}\Big)(f)\cdot\overline{\mathcal{F}\Big(\mathfrak{S}^{(l-1)}\{\mathbf{v}\}\Big)}(f^{\prime})\bigg].$ (9)
$\widetilde{\mathcal{G}^{(l)}}\{\mathbf{u},\mathbf{v}\}(f,f^{\prime}):=\delta(f-f^{\prime})\frac{C_{R^{(l)}}}{n^{(l-1)}}\sum_{j}\mathcal{F}\Big(\mathfrak{S}^{(l-1)}\{\mathbf{u}\}\Big)_{j}(f)\overline{\mathcal{F}\Big(\mathfrak{S}^{(l-1)}\{\mathbf{v}\}\Big)_{j}}(f^{\prime}).$ (10)
@@ 3.2 Running of Couplings
In this subsection we derive recursion relations for the correlation functions—what, in field theory language, is the running of couplings. As the layer width tends to infinity, the pre-activation ensemble becomes Gaussian and the layer metric converges to a deterministic kernel $\mathcal{K}^{(l)}$ (see Appendix B). In this vanishing fluctuation regime, the next layer kernel can be expressed as a Gaussian expectation with covariance $\mathcal{K}^{(l)}$ inherited from the previous layer; equivalently, $\mathcal{K}^{(l+1)}$ is obtained by integrating over the Gaussian process defined by $\mathcal{K}^{(l)}$:
$\begin{split}\mathcal{K}^{(l+1)}\{\mathbf{u},\mathbf{v}\}(f,f^{\prime}):=\delta(f-f^{\prime})C_{R^{(l+1)}}(f)\langle\mathcal{F}(\mathfrak{S}^{(l)}\{\mathbf{u}\}),\mathcal{F}(\mathfrak{S}^{(l)}\{\mathbf{v}\})\rangle_{\mathcal{K}^{(l)}}(f,f^{\prime}).\end{split}$
@@
$\begin{split}\mathcal{K}^{(1)}\{\mathbf{u},\mathbf{v}\}(f,f^{\prime}):=\frac{C_{W}}{n^{(0)}}\sum_{j}\hat{u_{j}}(f)\hat{v_{j}}(f^{\prime}).\end{split}$ (11)
$\begin{split}\mathcal{K}^{(l+1)}\{\mathbf{u}\}(f,f^{\prime}):=\delta(f-f^{\prime})\frac{C_{R^{(l+1)}}}{n^{(l)}}g(\mathcal{K}^{(l)}),\\ g(\mathcal{K}):=\langle\mathcal{F}(\mathfrak{S}^{(l)}\{\mathbf{u}\}),\mathcal{F}(\mathfrak{S}^{(l)}\{\mathbf{u}\})\rangle_{\mathcal{K}^{(l)}}(f,f^{\prime}).\end{split}$ (12)
@@
$\tilde{\chi}_{\perp}(f,f^{\prime}):=\|\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\})\|_{\mathcal{K}^{(l)}\{u_{0}\}}(f,f^{\prime})$
@@
Here, $\tilde{\chi}_{\perp}$ is calculated at layer $l$ and $C(f)$ is $C_{R^{(l+1)}}(f)$.
4 Results: Three Representative Cases
In this section we put the general framework of Section 3 to work by carrying out the calculation in three representative settings—(i) analytic activations, (ii) scale-invariant architectures, and (iii) networks equipped with residual connections.